OpenStudy (anonymous):

In a trial of 136 patients, 34 claimed to experience head aches from a test medicine. a) what is the 95% confidence interval in ascending order? b)Interpret the confidence interval.

OpenStudy (anonymous):

You have an estimated proportion of \(\hat{p}=\dfrac{34}{136}=0.25\). The \((1-\alpha)\times100\%\) confidence interval for a proportion is given by \[\left(\hat{p}-Z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}},~\hat{p}+Z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\right)\] where \(Z_{\alpha/2}=Z_{0.025}=1.96\). The sample size is \(n=136\).